OpenStudy (shamil98):

Each molecule of hydrogen gas (H2) is made of two hydrogen atoms. If you have 6.022 × 1023 molecules of hydrogen gas, then how many moles of hydrogen atoms do you have? 1 (its one right? ) 2 6 12

OpenStudy (abb0t):

6.022 \(\times\) 10\(^{23}\) molecules \(\times\) \(\frac{mol}{6.022 \times 10^{23} ~atom}\)

OpenStudy (abb0t):

oops, bottom is supposed to be molecule, not atom.